# Conformal Validity Guarantees Exist for Any Data Distribution

### v1

@@ Abstract
As machine learning (ML) gains widespread adoption, practitioners are increasingly seeking means to quantify and control the risk these systems incur. This challenge is especially salient when ML systems have autonomy to collect their own data, such as in black-box optimization and active learning, where their actions induce sequential feedback-loop shifts in the data distribution. Conformal prediction has emerged as a promising approach to uncertainty and risk quantification, but existing variants either fail to accommodate sequences of data-dependent shifts, or do not fully exploit the fact that agent-induced shift is under our control. In this work we prove that conformal prediction can theoretically be extended to any joint data distribution, not just exchangeable or quasi-exchangeable ones, although it is exceedingly impractical to compute in the most general case. For practical applications, we outline a procedure for deriving specific conformal algorithms for any data distribution, and we use this procedure to derive tractable algorithms for a series of agent-induced covariate shifts. We evaluate the proposed algorithms empirically on synthetic black-box optimization and active learning tasks.
@@ 2.1 Standard Conformal Prediction
$\mathbb{P}[Y_{n+1}\in\widehat{\mathcal{C}}_{n}(X_{n+1})]\geq 1-\alpha.$ (1)
@@
Standard CP methods assume that all the data $Z_{1},\dots,Z_{n+1}$ are exchangeable—that is, that their joint distribution is invariant to permutations, a special case being independent and identically distributed (IID). Under exchangeability, $\widehat{\mathcal{C}}_{n}(x)$ in Eq. (2) achieves the valid coverage guarantee in Eq. (1). This standard CP coverage guarantee is distribution-free in the sense of making no assumptions about the specific parametric form of the joint data distribution—this is not to be confused with robustness to distribution shift, which violates exchangeability for example when the distribution of $Z_{n+1}$ differs from that of $Z_{1:n}$. Exchangeability is thus still a strong assumption in its own way. For instance, it excludes many environment-dependent or time-dependent distribution shifts which often occur in practical settings.
@@ 2.3 Multistep Feedback Covariate Shift
Whereas standard and feedback covariate shift are “one-step” distribution shifts between training and test data batches, in this work we focus our practical implementations on a more general setting we call multistep feedback covariate shift (MFCS). MFCS is not only a clear example of a sequentially-dependent data distribution that goes far beyond exchangeability, but it moreover characterizes important instances of feedback-loop shifts inherent to the actions of AI/ML agents. For each time $t\in\{1,...,T\}$, MFCS allows $X_{n+t}$ to change arbitrarily depending on the past data $Z_{1:(n+t-1)}=\{Z_{i}\}_{i=1}^{n+t-1}$, while assuming that $Y\mid X$ remains invariant. More precisely, MFCS assumes
@@
Note that $n$ is the initial sample size of an IID (or exchangeable, more generally) initialization, while $t$ is the number of non-exchangeable MFCS steps past this initialization. Setting $T=1$ immediately reduces Eq. (3) to one-step FCS as considered by Fannjiang et al. (2022) and moreover to standard covariate shift if independence is further assumed.
@@ 3 Related Work
Conformal prediction under distribution shift. Prior work can broadly be categorized into variants of weighted conformal prediction (Tibshirani et al., 2019; Podkopaev & Ramdas, 2021; Fannjiang et al., 2022; Prinster et al., 2022, 2023; Barber et al., 2023; Farinhas et al., 2023) and adaptive conformal prediction (Gibbs & Candes, 2021; Zaffran et al., 2022; Angelopoulos et al., 2023; Feldman et al., 2023; Bhatnagar et al., 2023). The former approach exploits knowledge or estimates of the distribution shift to proactively adjust inference and thereby achieve guarantees that hold at inference time; meanwhile, the latter aims to maintain a target risk threshold by retroactively responding to threshold violations and thereby obtain guarantees most meaningful “in the long run” rather than at a given inference time.
This work builds on the literature of weighted CP, and our practical applications focus on cases of MFCS where the shift is agent-induced and thus known. Aside from Tibshirani et al. (2019) and Fannjiang et al. (2022), Prinster et al. (2022) and Prinster et al. (2023) also studied CP under (one-step) standard or feedback covariate shift by extending cross-validation-style CP methods (from Barber et al. (2021)) to those settings for flexibly balancing compute-versus-sample efficiency. Otherwise, our work is most closely related to Barber et al. (2023), which bounds the worst-case “coverage gap” (i.e., loss of coverage below Eq. 1) if the conformal weights are unknown and instead fixed a priori independently of the data. In contrast, our extensions of CP to non-exchangeable distributions use data-dependent weights to achieve coverage validity guarantees as in Eq. 1 (without a gap in coverage).
Conformal inference for decision-making. There is a growing body of work applying the ideas of CP to decision-making to inform utility estimates (Vovk & Bendtsen, 2018; Stanton et al., 2023; Salinas et al., 2023) or satisfy constraints (Zhang et al., 2023; Dixit et al., 2023; Laufer-Goldshtein et al., 2023; Jin & Candès, 2023). In particular, Stanton et al. (2023) drew motivation from Fannjiang et al. (2022) to combine CP with Bayesian optimization but noted that a gap in theory prevented a formal coverage guarantee after the first round, a gap we address in this work.
4 Theoretical Contributions and Methods
@@ 4.1 The Role of (Weighted) Exchangeability and Related Assumptions in Conformal Prediction
Informal Statement of the Key Insight One can think of conformal prediction (CP) as an inverted permutation test (with a conservative adjustment). Informally, our first key insight is that—from this perspective of an inverted permutation test—the only role of exchangeability, weighted exchangeability (Tibshirani et al., 2019), or related distributional assumptions such as pseudo-exchangeability (Fannjiang et al., 2022) in CP is to enable tractability by telling us how much to “weigh” or “count” a given permutation of datapoints (or score values); in particular, by facilitating computation of the relative likelihood of each sequence of observations, relative to some core function (that does not depend on the ordering if its inputs), given the unordered “bag” of values $\{z_{1},...,z_{n+1}\}$ (see Appendix C details on these definitions and their roles). So, extensions of CP should be possible so long as they are able to quantify the relative likelihood of any ordering of datapoints (or scores).
Formal Sketch of the Key Insight We now sketch with formal notation our main insight. This sketch will correspond to viewing CP as an inverted permutation test, provide intuition for our main theorem in Section 4.2, and moreover highlight key ideas underlying the theorem’s first proof step.
Constructing CP from an inverted permutation test begins by conditioning on the set [^4] For simplicity, we assume all the values are distinct to work with sets rather than multisets; in general, the argument holds with more complex notation using uniform randomness to break ties. of data values $\{z_{1},...,z_{n+1}\}$—that is, let $E_{z}$ denote the event that $\{Z_{1},...,Z_{n+1}\}=\{z_{1},...,z_{n+1}\}$. To be clear, this notation means that the unordered “bag” of values is known but not whether $Z_{i}=z_{i}$, and so on; that is, we know the random variable $Z_{i}$ takes on an observed value in $\{z_{1},...,z_{n+1}\}$, but we do not yet know which value. Conditioning on $E_{z}$, we are then interested in the probability that the test point’s score $V_{n+1}$ was actually $v_{i}$, for each $i\in\{1,...,n+1\}$, which we can write as
@@
where $f$ denotes the joint probability density function (PDF). Standard CP proceeds directly from this statement by leveraging exchangeability to show that every permutation of the data values is equally likely, thus reducing the problem to counting permutations; weighted CP proceeds by factorizing $f$ using weighted exchangeability, with the product of weight functions $\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})$ telling us how much weight to give to the permutation $\sigma$; and, CP with pseudo-exchangeable data proceeds by factorizing $f$ similarly, with more explicit weight functions describing a specific instance of potential dependencies (See Appendix C.3).
More generally, the key insight comes from observing that equation (4) does not rely on any assumptions about the joint distribution $f$. This can be seen by deriving the statement using Bayes rule and the law of total probability (LOTP):
@@
That is, the last step uses the LOTP in both the numerator and the denominator—for instance, in the denominator the summation is over all the possible ways that the event $E_{z}$ could occur, which amounts to all the different possible permutations $\sigma$ of the values. The crux of extending CP to any, possibly non-exchangeable joint distribution can thus reduce to the problem of either computing $f(z_{\sigma(1)},...,z_{\sigma(n+1)})$ directly for any ordering $\sigma$ of the observations, or to be able to factorize $f$ into terms that are tractable to compute (e.g., weight or factor functions) and remaining terms that do not depend on $\sigma$ (e.g., a core measure that is invariant to permutations and thus cancels out in the ratio).
@@ 4.2 A General View of Conformal Prediction
Main Formal Result With the insight from Section 4.1 in mind, we are now ready to see how—in theory—extensions of CP are possible for any, potentially non-exchangeable distribution [^5] To be precise, we require a non-singular distribution so that there exists a valid PDF $f$, or more generally, a valid Radon-Nikodym derivative with respect to an arbitrary base measure. of the calibration and test data $Z_{1},...,Z_{n+1}$ (we let $t=1$ for simplicity, but in general our argument holds for $Z_{1},...,Z_{n+t}$). This section’s theorem formally states this result as the coverage guarantee for a generalized weighted CP set for any joint probability density function (PDF) $f$. We will then soon discuss both the practical benefits and limitations of such a general view.
@@
Further Discussion of Tibshirani et al. (2019) This general view of CP is implicit in a close reading of Tibshirani et al. (2019), however the authors did not explicitly state or prove the general result we present in Theorem 4.1. Instead, Tibshirani et al. (2019) premised their main result on their definition of weighted exchangeability despite this condition not being strictly necessary for CP in the broad sense. While we maintain that weighted exchangeability is still a useful definition (see Appendix C.2), subsequent literature has seemingly been led to believe that new definitions of “quasi-exchangeability” are necessary for advancing and delineating the boundaries of CP theory—for example, Fannjiang et al. (2022)’s introduction of pseudo-exchangeability to study the single-step FCS setting. We hope our contribution will lead to wider recognition of the generality of Tibshirani et al. (2019)’s analysis and foster a broadened scope of possibility in the study of CP: in particular, we encourage a shift away from viewing “quasi-exchangeability” conditions as theoretical boundaries to be overcome, and instead towards primarily using such conditions to clearly specify the assumptions granted by a given setting of interest or adopted for practical tractability.
4.3 Deriving Weighted CP Algorithms for MFCS, or Any Data Distribution
A General Procedure for Deriving Weighted CP Algorithms with Valid Coverage Guarantees To improve the accessibility and practical utility of our main theoretical contribution, we now outline a general procedure that can be used to derive weighted CP algorithms for any specific joint PDF $f$ over the calibration and test data. Theorem 4.1 implies the existence of this procedure without premising it on any assumptions on $f$, while emphasizing that the only role of such assumptions is for tractability in the last step.
@@ 4.3 Deriving Weighted CP Algorithms for MFCS, or Any Data Distribution
    Factorize $f$: Factorize the joint PDF $f$ using standard probability rules (e.g., Bayes or chain rule) and assumptions from Step 1 to separate dynamic factors from those that are invariant (to permutations $\sigma$ of the data indices). (For example, for MFCS $f$ can be written as a product of dynamic factors $X_{j}\mid Z_{1},…,Z_{j-1}$ and invariant factors $Y\mid X$.)
@@
Deriving CP for Agent-Induced MFCS For a concrete demonstration of how this general procedure can be used to derive novel CP methods, we now return to the MFCS setting introduced in Section 2.3, which characterizes common instances of agent-induced feedback-loop data shifts. We provide only a sketch of key ideas here; for full details see Appendix B. Step 1 of this procedure for MFCS begins by recalling Eq. (3), the conditional independence or invariance assumptions that we used to define MFCS. Using standard probability rules and leveraging these MFCS assumptions, for Step 2 we are able to factorize the joint PDF $f$ to separate the time-dependent factors corresponding to $X_{j}\mid Z_{1},…,Z_{j-1}$ from the time-invariant factors corresponding to $Y\mid X$:
@@ 5 Experimental Results
We evaluate the practical performance of our proposed CP methods for ML-agent-induced MFCS on synthetic black-box optimization (specifically protein design) and active learning tasks. All experiments focus on a regression setting with all CP methods using the absolute residual score $\widehat{\mathcal{S}}(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}_{t}(x)$ is the predicted mean. Our proposed methods are defined by estimating Eq. (6) with weights as in Eq. (9). The primary desired criterion for evaluating a method’s reliability is whether its prediction intervals maintain empirical coverage at or above the target coverage level $1-\alpha$. Secondarily—if target coverage is achieved—then smaller predictive sets are more informative (i.e., predicting the full label space $\widehat{\mathcal{C}}_{n}(x)=\mathcal{Y}$ could trigger deferral to an expert but does not itself convey new insights from the predictor). In all experiments we compare to Exchangeable CP (Vovk et al., 2005) and One-Step FCS CP (Fannjiang et al., 2022), and in the active learning experiments we additionally compare with Adaptive Conformal Inference (ACI) (Gibbs & Candes, 2021). Our proposal distributions follow the form $\mathbb{P}_{t}(X_{n+t}=x):=\mathbb{P}(X_{n+t}=x\mid Z_{1:(n+t-1)})\propto\exp(\lambda\cdot u_{t}(x))$, where $u_{t}$ is some utility function and $\lambda\geq 0$ is a hyperparameter that controls the magnitude of agent-induced covariate shift. In all experiments we corrupt observed labels with Gaussian noise to simulate noisy observations. We add every query point to the training set for the full CP experiments, and for split CP we randomly assign the query points to either training or calibration by flipping a fair coin. (Details in Appendix D.)

### v2

@@ Abstract
As machine learning (ML) gains widespread adoption, practitioners are increasingly seeking means to quantify and control the risk these systems incur. This challenge is especially salient when ML systems have autonomy to collect their own data, such as in black-box optimization and active learning, where their actions induce sequential feedback-loop shifts in the data distribution. Conformal prediction has emerged as a promising approach to uncertainty and risk quantification, but prior variants’ validity guarantees have assumed some form of “quasi-exchangeability” on the data distribution, thereby excluding many types of sequential shifts. In this paper we prove that conformal prediction can theoretically be extended to any joint data distribution, not just exchangeable or quasi-exchangeable ones, although it is exceedingly impractical to compute in the most general case. For practical applications, we outline a procedure for deriving specific conformal algorithms for any data distribution, and we use this procedure to derive tractable algorithms for a series of ML-agent-induced covariate shifts. We evaluate the proposed algorithms empirically on synthetic black-box optimization and active learning tasks.
@@ 2.1 Standard Conformal Prediction
$\mathbb{P}\big{\{}Y_{n+1}\in\widehat{\mathcal{C}}_{n}(X_{n+1})\big{\}}\geq 1-\alpha.$ (1)
@@
Standard CP methods assume that all the data $Z_{1},\dots,Z_{n+1}$ are exchangeable—that is, that their joint distribution is invariant to permutations, a special case being independent and identically distributed (IID). Under exchangeability, $\widehat{\mathcal{C}}_{n}(x)$ in Eq. (2) achieves the valid coverage guarantee in Eq. (1) with finite samples. This standard CP guarantee is distribution-free in the sense of making no assumptions about the specific parametric form of the data distribution—this is not to be confused with robustness to distribution shift, which violates exchangeability for example when the distribution of $Z_{n+1}$ differs from that of $Z_{1:n}$. Exchangeability is thus still a strong assumption in its own way. For instance, it excludes many environment-dependent or time-dependent distribution shifts which often occur in practice.
@@ 2.3 Multistep Feedback Covariate Shift
Whereas standard and feedback covariate shift are “one-step” shifts between training and test data batches, in this work we focus our practical evaluations on a more general setting we call multistep feedback covariate shift (MFCS). MFCS is not only a clear example of a sequentially-dependent data distribution that goes far beyond exchangeability, but it moreover describes important instances of feedback-loop shifts inherent to the actions of AI/ML agents. For each time $t\in\{1,...,T\}$, MFCS allows $X_{n+t}$ to change depending on the past data $Z_{1:(n+t-1)}=\{Z_{i}\}_{i=1}^{n+t-1}$, while assuming that $Y|X$ remains invariant. More precisely, MFCS assumes
@@
Note that $n$ is the initial sample size of an IID initialization, and $t$ is the number of MFCS steps past this initialization. Setting $T=1$ reduces Eq. (3) to one-step FCS (Fannjiang et al., 2022) and moreover to standard covariate shift if independence is further assumed; we also note that MFCS is similar to “sequential ignorability” in causal inference.
@@ 3 Related Work
Conformal prediction under distribution shift. Prior work can largely be categorized into variants of weighted CP (Tibshirani et al., 2019; Podkopaev & Ramdas, 2021; Xu & Xie, 2021; Fannjiang et al., 2022; Prinster et al., 2022, 2023; Barber et al., 2023; Farinhas et al., 2023; Nair & Janson, 2023; Yang et al., 2024) and adaptive CP (Gibbs & Candès, 2021, 2022; Zaffran et al., 2022; Angelopoulos et al., 2023; Feldman et al., 2023; Bhatnagar et al., 2023). The former exploits knowledge or estimates of the data shift to proactively adjust inference and thereby achieve guarantees that hold at inference time; meanwhile, the latter aims to maintain a target risk threshold by retroactively responding to threshold violations and thus obtain guarantees most meaningful “in the long run” rather than at inference time.
This work builds on the weighted CP literature, and our practical applications focus on cases of MFCS where the shift is agent-induced and thus known. Aside from Tibshirani et al. (2019) and Fannjiang et al. (2022), Prinster et al. (2022) and Prinster et al. (2023) also studied CP under (one-step) standard or feedback covariate shift by extending cross-validation-style CP methods (from Barber et al. (2021)) to those settings for flexibly balancing compute-versus-sample efficiency. Nair & Janson (2023) studied a “$Y$-stationarity” setting similar to MFCS with a Monte Carlo-based CP algorithm for discrete data. Otherwise, our work relates to Barber et al. (2023), which bounds the worst-case “coverage gap” (loss of coverage below Eq. (1)) if the CP weights are unknown and instead fixed a priori independently of the data. In contrast, we study CP for non-exchangeable distributions using data-dependent weights to achieve coverage validity guarantees as in Eq. (1) (without a coverage gap).
Conformal inference for decision-making. Ideas from CP are increasingly being applied to decision-making to inform utility estimates (Vovk & Bendtsen, 2018; Stanton et al., 2023; Salinas et al., 2023; Lekeufack et al., 2023) or satisfy constraints (Zhang et al., 2023; Dixit et al., 2023; Laufer-Goldshtein et al., 2023; Jin & Candès, 2023). In particular, Stanton et al. (2023) drew motivation from Fannjiang et al. (2022) to combine CP with Bayesian optimization but noted that a gap in theory prevented a formal coverage guarantee after the first round, a gap we address in this work.
4 Theory and Method Contributions
@@ 4.1 The Role of (Weighted) Exchangeability and Related Assumptions in Conformal Prediction
Informal Statement of the Key Insight One can think of CP as an inverted permutation test (with a conservative adjustment). Informally, the key insight is that—from this perspective of an inverted permutation test—the only role of exchangeability, weighted exchangeability (Tibshirani et al., 2019), or related distributional assumptions such as pseudo-exchangeability (Fannjiang et al., 2022) in CP is to enable tractability by telling us how much to “weigh” or “count” a given ordering of datapoints; in particular, by facilitating computation of the relative likelihood of each sequence of observations, relative to some (permutation-invariant) core function, given the unordered “bag” of observed values (details in Appendix C). So, extensions of CP should be possible so long as one can quantify the probability of any ordering of observed datapoints (or score values).
Formal Sketch of the Key Insight We first sketch analysis from Tibshirani et al. (2019) (which implicitly constructs CP as an inverted permutation test) followed by the key insight for our current paper’s main result. Their analysis begins by conditioning on the set [^4] For simplicity, we assume all the values are distinct to work with sets rather than multisets; in general, the argument holds with more complex notation using uniform randomness to break ties. of data values (or scores)—that is, on the event $\{Z_{1},...,Z_{n+1}\}=\{z_{1},...,z_{n+1}\}$ denoted by $E_{z}$. To be clear, $E_{z}$ means the unordered “bag” of values is known but not whether $Z_{i}=z_{i}$, and so on; we know that the random variable $Z_{i}$ takes on an observed value in $\{z_{1},...,z_{n+1}\}$, but we do not yet know which value (similarly for the scores $V_{i}$). Given $E_{z}$, the goal is to examine the probability that the test point’s score $V_{n+1}$ was actually $v_{i}$, for each $i\in\{1,...,n+1\}$, which they write as
@@
where $f$ denotes the joint probability density function (PDF). Standard CP proceeds from this statement by substituting $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f(z_{1},...,z_{n+1})$ via exchangeability, reducing the problem to counting permutations; weighted CP proceeds by factorizing $f$ using weighted exchangeability, with the product of weight functions representing the “weight” given to each permutation $\sigma$; and, CP with pseudo-exchangeable data proceeds similarly to handle a specific instance of potential dependencies (details in Appendix C).
Generally, the key insight in our paper is that simplifying Eq. (4) is only a practical rather than a theoretically necessary step. For intuition on this point, note that Eq. (4) makes no assumptions on $f$; it follows from the definition of conditional probability and law of total probability (LOTP):
@@
That is, the last step uses LOTP in both the numerator and denominator—for instance, the denominator’s sum is over every way the event $E_{z}$ could occur, which amounts to all the possible permutations $\sigma$ of the values. The crux of extending CP to any, possibly non-exchangeable $f$ can thus reduce to computing $f(z_{\sigma(1)},...,z_{\sigma(n+1)})$ directly for all orderings $\sigma$, or factoring $f$ into “dyanmic” terms that are tractable to compute (e.g., “weight functions”) and remaining permutation-invariant terms that cancel in the ratio.
@@ 4.2 A General View of Conformal Prediction
Main Formal Result With the insight from Section 4.1 in mind, we are now ready to see how—in theory—extensions of CP are possible for any, potentially non-exchangeable distribution of the calibration and test data $Z_{1},...,Z_{n+1}$ with a valid PDF [^5] To be precise, we require a non-singular distribution so that there exists a valid PDF $f$, or more generally, a valid Radon-Nikodym derivative with respect to an arbitrary base measure. $f$ (we let $t=1$ for simplicity, but in general our argument holds for $Z_{1},...,Z_{n+t}$). This section’s theorem formally states this result as the coverage guarantee for a generalized weighted CP set for arbitrary $f$.
@@
Further Discussion of Tibshirani et al. (2019) This general view of CP is implicit in a close reading of Tibshirani et al. (2019), [^6] We note that the main insight in our paper is also invoked in the following lecture notes (Tibshirani, 2023): https://www.stat.berkeley.edu/~ryantibs/statlearn-s23/lectures/conformal_ds.pdf, which a sharp reader could use to infer Theorem 4.1. However, based on how Tibshirani et al. (2019) has been cited in the CP literature (i.e., primarily as extending CP to standard covariate shift or to weighted exchangeability), we believe this important insight has not been fully appreciated by the community and we hope that our contribution can help correct this oversight. but in that paper the authors did not explicitly state or prove the general result we present in Theorem 4.1. Instead, Tibshirani et al. (2019) premised their main result on their definition of weighted exchangeability, although this condition is not strictly necessary for CP in the broad sense. While we maintain that weighted exchangeability is still a useful definition (see Appendix C.2), subsequent literature has seemingly been led to believe that new definitions of “quasi-exchangeability” are necessary for advancing and delineating the boundaries of CP theory—for example, Fannjiang et al. (2022)’s introduction of pseudo-exchangeability to study the single-step FCS setting. We hope our contribution will lead to wider recognition of the generality of Tibshirani et al. (2019)’s analysis and foster a broadened scope of possibility in the study of CP: in particular, we encourage a shift away from viewing “quasi-exchangeability” conditions as theoretical boundaries to be overcome, and instead towards primarily using such conditions to clearly specify the assumptions granted by a given setting of interest or adopted for practical tractability.
4.3 Deriving Weighted CP Validity Guarantees and Algorithms for MFCS, or Any Data Distribution
A General Procedure for Deriving Weighted CP Algorithms with Valid Coverage Guarantees To improve the accessibility and practical utility of our paper’s main theoretical contribution, we now outline a general procedure that can be used to derive weighted CP algorithms for any specific joint PDF $f$ over the calibration and test data. Theorem 4.1 implies the existence of this procedure without premising it on any assumptions on $f$, while emphasizing that the only role of distributional assumptions like exchangeability is for tractability in the last step.
@@ 4.3 Deriving Weighted CP Validity Guarantees and Algorithms for MFCS, or Any Data Distribution
    Factorize $f$: Factorize the joint PDF $f$ using standard probability rules (e.g., conditional probability, chain rule, etc.) and assumptions from Step 1 to separate dynamic factors from those that are invariant (to permutations $\sigma$ of the data indices). (For example, for MFCS $f$ can be written as a product of dynamic factors $X_{j}\mid Z_{1},…,Z_{j-1}$ and invariant factors $Y\mid X$.)
@@
Deriving CP for Agent-Induced MFCS For a concrete demonstration of how this general procedure can be used to derive CP methods for a given practical setting, we now return to the MFCS setting introduced in Section 2.3, which characterizes common instances of ML-agent-induced feedback-loop data shifts. We provide only a sketch of key ideas here; for full details see Appendix B. Step 1 of this procedure for MFCS begins by recalling Eq. (3), the conditional independence or invariance assumptions that we used to define MFCS. Using standard probability rules and leveraging these MFCS assumptions, for Step 2 we are able to factorize the joint PDF $f$ to separate the time-dependent factors corresponding to $X_{j}\mid Z_{1},…,Z_{j-1}$ from the time-invariant factors corresponding to $Y\mid X$:
@@ 5 Experimental Results
We evaluate the practical performance of our proposed CP methods for ML-agent-induced MFCS on synthetic black-box optimization (specifically protein design) and active learning tasks. All experiments focus on a regression setting with all CP methods using the absolute residual score $\widehat{\mathcal{S}}(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}_{t}(x)$ is the predicted mean. Our proposed methods are defined by estimating Eq. (6) with weights as in Eq. (9). The primary desired criterion for evaluating a method’s reliability is whether its prediction intervals maintain empirical coverage at or above the target coverage level $1-\alpha$. Secondarily—if target coverage is achieved—then smaller predictive sets are more informative (i.e., predicting the full label space $\widehat{\mathcal{C}}_{n}(x)=\mathcal{Y}$ could trigger deferral to an expert but does not itself convey new insights from the predictor). In all experiments we compare to Exchangeable CP (Vovk et al., 2005) and One-Step FCS CP (Fannjiang et al., 2022), and in the active learning experiments we additionally compare with Adaptive Conformal Inference (ACI) (Gibbs & Candès, 2021). Our proposal distributions follow the form $\mathbb{P}_{t}(X_{n+t}=x):=\mathbb{P}(X_{n+t}=x\mid Z_{1:(n+t-1)})\propto\exp(\lambda\cdot u_{t}(x))$, where $u_{t}$ is some utility function and $\lambda\geq 0$ is a hyperparameter that controls the magnitude of agent-induced covariate shift. In all experiments we corrupt observed labels with Gaussian noise to simulate noisy observations. We add every query point to the training set for the full CP experiments, and for split CP we randomly assign the query points to either training or calibration by flipping a fair coin. (Details in Appendix D.)